# Linear constellations in primes with arithmetic restrictions

### v2

@@ Abstract.
We prove analogues of the Green-Tao-Ziegler theorem on linear constellations in primes, in which the primes under consideration are restricted by certain arithmetic conditions. Our first main result is conditional upon Hooley’s Riemann hypothesis and imposes the extra condition that the primes have prescribed primitive roots. Our second main result is unconditional and imposes the extra condition that the primes have prescribed Artin symbols in given Galois number fields. In the appendix we present an application of the second result in inverse Galois theory.
@@ 1.1. Primes with prescribed primitive roots
simultaneously take values in the set of primes with primitive root $5$. This is a constellation of complexity $2$ in the sense of Green and Tao [13]. Their celebrated result together with Ziegler [17] gives asymptotics for all constellations of finite complexity in unrestricted primes, implying in particular the existence of infinitely many solutions $(n_{1},n_{2})$ to our question if one ignores the primitive root condition.
@@
Here, $\mu_{k}$ denotes the group of $k$-th roots of unity in $\mathbb{C}$. For background and history surrounding Artin’s conjecture and Hooley’s result, we recommend Moree’s survey [32]. Our first main result establishes, conditionally on $\operatorname{HRH}(7)$, the existence of infinitely many solutions $(n_{1},n_{2})$ as above. In fact, we allow arbitrary primitive roots, arbitrary finite-complexity systems of affine-linear forms, and obtain an asymptotic formula analogous to the Green-Tao-Ziegler theorem for unrestricted primes.
@@
#$\#\left\{p\leq N\ :\ p\equiv b\bmod q,\ \mathbb{F}_{p}^{\times}=\langle a\rangle\right\}=\delta(a,b,q)\frac{N}{\log N}+o_{a,q}\left(\frac{N}{\log N}\right).$ (1.2)
@@
$\Lambda_{a}(n):=\begin{cases}\log p,&\text{ if }n=p^{e}\text{ with $e\in\mathbb{N}$,\ $p$ prime and }\mathbb{F}_{p}^{\times}=\langle a\rangle,\\ 0,&\text{ otherwise},\end{cases}$ (1.5)
@@
where we have used the notation $[N]=\{1,\ldots,N\}$. Our main result extends this from $\psi(n)=b+qn$ to arbitrary finite complexity [^1] Recall the characterisation of finite complexity from [13, Lemma 1.6]: no two of the forms $\psi_{i}$ have linearly dependent linear parts. systems of affine-linear forms $\Psi=(\psi_{1},\ldots,\psi_{t}):\mathbb{Z}^{s}\to\mathbb{Z}^{t}$ as in [13, Definition 1.1 and Definition 1.5]. For $N\in\mathbb{N}$, the size of $\Psi$ relative to $N$ is defined as
@@
where $\dot{\psi}_{i}=\psi_{i}-\psi_{i}(0)$ is the linear part of $\psi_{i}$ and $e_{1},\ldots,e_{s}$ is the standard basis of $\mathbb{Z}^{s}$. Let $\mathbf{a}=(a_{1},\ldots,a_{t})\in\mathbb{Z}^{t}$ such that no $a_{i}$ is equal to $-1$ or a perfect square. For any $q\in\mathbb{N}$, we define the density
@@
[…]
With this setup in place, we can state our first main result, a version of [13, Main Theorem] with prescribed primitive roots, conditional upon the minimal version of GRH that is currently needed to even establish the existence of infinitely many primes with these prescribed primitive roots.
@@
Let $s,t,L,N\in\mathbb{N}$. Let $\mathbf{a}=(a_{1},\ldots,a_{t})\in\mathbb{Z}^{t}$ such that no $a_{i}$ is equal to $-1$ or a perfect square, and assume that $\operatorname{HRH}(a_{i})$ holds for all $1\leq i\leq t$. Let $\Psi=(\psi_{1},\ldots,\psi_{t}):\mathbb{Z}^{s}\to\mathbb{Z}^{t}$ be a system of affine-linear forms of finite complexity with size $\lVert\Psi\rVert_{N}\leq L$. Let $X\subseteq[-N,N]^{s}$ be a convex set. Then
@@
for unrestricted primes in [13, (1.6)], where $\phi(q)$ denotes Euler’s totient and for any proposition $P$ we write $\mathbf{1}_{P}=1$ if $P$ is true and $\mathbf{1}_{P}=0$ otherwise. Here, $\mathbf{1}_{\gcd(\psi_{i}(n),q)=1}/\phi(q)$ can be seen as the probability that a random prime is congruent to $\psi_{i}(n)\bmod q$ by the prime number theorem in arithmetic progressions. Contrary to [13], our factor $\mathfrak{S}(\Psi,\mathbf{a})$ does not factor completely as an Euler product, hence the need to introduce the quantity $\mathscr{D}_{\mathbf{a}}$. Moreover, our formula features an additional global factor $\prod_{i=1}^{t}\delta(a_{i},0,1)$, where of course $\delta(a_{i},0,1)$ should be seen as the probability that a random prime has primitive root $a_{i}$. Both of these phenomena were already observed in a special case in [10], see Example 1.8 below.
@@
If $\sigma_{\mathbf{a},\Psi}(q)=0$ for some $q\in\{\mathscr{D}_{\mathbf{a}}\}\cup\{p\ :\ p\nmid\mathscr{D}_{\mathbf{a}},\ p\ll_{s,t,L,\mathbf{a}}1\}$, then for each $c\in(\mathbb{Z}/q\mathbb{Z})^{s}$ there is some $i\in[t]$ with $\delta(a_{i},\psi_{i}(c),q)=0$. From Moree’s explicit description of $\delta(a,b,q)$, see §3.3, it is straightforward to see that then every prime congruent to $\psi_{i}(c)\bmod q$ with primitive root $a_{i}$ must lie in the set $\{2,\gcd(\psi_{i}(c),q)\}$. Taking all possible candidates for all $c$, we end up with a finite collection of cases, in each of which one of the forms $\psi_{i}(n)$ has to take a fixed value. In each case, one can use linear algebra to either show that no solutions exist, or express the system in $s-1$ variables. However, this new system might have infinite complexity, so Theorem 1.2 might not apply any longer. See the end of §11.3 for an example.
@@ 1.2. Primes with prescribed Artin symbols
Our second main result is unconditional. Here, instead of prescribed primitive roots, we require the primes to have prescribed Artin symbols $[K/\mathbb{Q},p]$ for some finite Galois extensions $K/\mathbb{Q}$. Recall that this Artin symbol is defined for primes $p$ unramified in $K$ and yields a conjugacy class in the Galois group $\operatorname{Gal}(K/\mathbb{Q})$, the class of Frobenius elements of prime ideals $\mathfrak{p}$ of the ring of integers of $K$ lying above $p$. For any such conjugacy class $C$, we define the Chebotarev-von Mangoldt function
$\Lambda_{K,C}(n):=\begin{cases}\log p,&\text{ if }n=p^{e}\text{ with $e\in\mathbb{N}$, $p$ prime, unramified in $K$ with }[K/\mathbb{Q},p]=C,\\ 0,&\text{ otherwise.}\end{cases}$ (1.9)
@@
$\eta_{K,C}(b,q):=\begin{cases}\frac{|C|}{[K(\mu_{q}):\mathbb{Q}]},&\begin{aligned} &\text{ if $\gcd(b,q)=1$ and $\sigma_{b}$ restricts to the same element}\\ &\text{ of $\operatorname{Gal}(K\cap\mathbb{Q}(\mu_{q})/\mathbb{Q})$ as the elements of $C$},\end{aligned}\\ 0,&\text{ otherwise.}\end{cases}$ (1.10)
@@
[…]
where $\Phi_{K_{i}^{\operatorname{ab}}}\in\mathbb{N}$ is the finite part of the conductor [^2] In concrete terms, $\Phi_{K_{i}^{\operatorname{ab}}}$ is the minimal $f\in\mathbb{N}$ such that $K_{i}^{\operatorname{ab}}\subseteq\mathbb{Q}(\mu_{f})$. Such $f$ exist by the Kronecker-Weber theorem. of the maximal abelian subextension $K_{i}^{\operatorname{ab}}/\mathbb{Q}$ of $K_{i}/\mathbb{Q}$. With this setup in place, our second main result is the following version of the Green-Tao-Ziegler theorem.
@@
The factor $\mathfrak{S}(\mathbf{K},\mathbf{C},\Psi)$ in the main term has an analogous interpretation as in Theorem 1.2. In particular, the quotient $\eta_{K_{i},C_{i}}(\psi_{i}(n),q)/\eta_{K_{i},C_{i}}(0,1)$ in (1.13) can be seen as the conditional probability that a random prime $p$ with Artin symbol $[K_{i}/\mathbb{Q},p]=C_{i}$ is congruent to $\psi_{i}(n)$ modulo $q$. In contrast to Theorem 1.2, these conditional probabilities simplify considerably: we will see in Lemma 3.4 that $\tau_{\mathbf{K},\mathbf{C},\Psi}(\mathscr{D}_{\mathbf{K}})=\beta_{\Psi,p}$ whenever $p\nmid\mathscr{D}_{K}$, with $\beta_{\Psi,p}$ the same density as for unrestricted primes in [13], defined in (1.8). Hence, we can also write
@@
which is very close to the formula for unrestricted primes in [13, Main Theorem].
@@
Analogous statements as in Remark 1.3 apply. One easily sees directly from the definition that if $\eta_{K,C}(b,q)=0$, then there is at most one possible prime $p\equiv b\bmod q$ with $[K/\mathbb{Q},p]=C$, namely $\gcd(b,q)$.
@@
The Green-Tao-Ziegler theorem [13] has been applied in combination with fibration or specialisation techniques in more algabraically-flavored contexts, where it is used to show that certain linear forms, e.g. the linear factors of a well-chosen discriminant polynomial, take prime values simultaneously infinitely often (see e.g. [18, 3, 24, 1]). Theorem 1.4 adds the capability of prescribing splitting conditions at these prime values, which we hope will be useful. To facilitate applications, we state the following implication of Theorem 1.4.
@@
Then there are infinitely many $n\in\mathbb{Z}^{s}\cap X$ for which the values $\psi_{i}(n)$ are distinct primes with $[K_{i}/\mathbb{Q},\psi_{i}(n)]=C_{i}$ for all $i$.
@@ 1.3. Application in inverse Galois theory
A special case of Corollary 1.6 was asked for explicitly by Kim and König in the context of inverse Galois problems with local conditions. In [24, Remark 6], they ask whether, for given $n\geq 2$, an analog of the Green-Tao-Ziegler theorem can prove the existence of infinitely many pairs $(s,t)\in\mathbb{Z}^{2}$, such that
[…]
@@
over $\mathbb{Q}$. In Appendix A, which is joint work with Joachim König, we apply Corollary 1.6 to answer this question positively for all $n\geq 2$. Using this to replace an application of the original Green-Tao-Ziegler theorem in the proof of [24, Theorem 5.5], we prove the following result on locally cyclic $\operatorname{S}_{n}$-extensions with prescribed Artin symbol at finitely many primes. This generalisation of the special case $n=5$ treated in [24, Theorem 5.5] was suggested by Kim and König as the motivation for their question. Recall that a Galois number field $K/\mathbb{Q}$ is locally cyclic if all decomposition groups are cyclic.
@@
    For each $p\in M_{2}$, we have $[K/\mathbb{Q},p]=C_{p}$.
This result is new even without the conditions (3) and (4). It answers positively a question of Bubboloni and Sonn [8], who asked whether locally cyclic Galois extensions of $\mathbb{Q}$ with Galois group $\operatorname{S}_{n}$ exist for all $n$. Let us note that [24, Theorem 5.5] also covers the case of $\operatorname{PGL}_{2}(7)$. A generalisation to $\operatorname{PGL}_{2}(p)$ for arbitrary primes $p$ follows from the main result of [1].
@@ 1.4. Previous results and examples
Actually, Kane’s formula is stated in terms of primes $p_{i}$ with $[K_{i}/\mathbb{Q},p_{i}]=C$ weighted by $\log(p_{i})$ instead of integers $n_{i}$ weighted by $\Lambda_{K_{i},C_{i}}(n_{i})$, but the contribution of prime powers making up the difference between these two settings is clearly negligible. Considering $c_{1},\ldots,c_{t}$ as fixed, we may assume that $M\ll N$, as otherwise the sum is zero. As explained in [13, Theorem 1.8] and its proof, one may choose a basis for the affine sublattice of $\mathbb{Z}^{t}$ given by $c_{1}x_{1}+\cdots+c_{t}x_{t}=M$, and thus describe it as the image of an affine-linear system $\Psi:\mathbb{Z}^{t-1}\to\mathbb{Z}^{t}$ of size $\lVert\Psi\rVert\ll 1$ and complexity $\leq 1$.
We describe Kane’s formula in §11.1 and show that it agrees with the corresponding special case of Theorem 1.4. In order to obtain his result, Kane generalises the proof of Vinogradov’s three primes theorem presented in [21, §19], based on classical analytic number theory surrounding the Hardy-Littlewood circle method. Some of Kane’s techniques play a role in our work. Most importantly, our Proposition 5.2 can be seen as a higher-order Fourier analysis version of [23, Lemma 21], Kane’s exponential sum estimate for polynomial phases evaluated at ideal norms.
We should mention that Kane’s result features a quantitative error term saving an arbitrary power of $\log N$, whereas our error terms are purely qualitative. It would be interesting to see whether the recent methods of Tao and Teräväinen [36], based on Manners’ quantitative inverse theorem for Gowers norms [28], by which they obtain a version of the Green-Tao-Ziegler theorem with quantitative error terms saving a small power of $\log\log N$, could be adapted to the situation of Theorem 1.4.
@@
of complexity $t-2$ takes infinitely many values in $\mathscr{P}^{t}$. Taking $\mathscr{P}$ to be the set of primes with $[K/\mathbb{Q},p]=C$ for some fixed $K,C$, which has positive density in the primes by (1.12) with $q=1$, one obtains a weakened special case of Theorem 1.4. Taking $\mathscr{P}$ to be the set of primes with primitive root $a$, one obtains a weakened special case of Theorem 1.2. In both cases, the precise asymptotic formulae in our results are new. In §11.3, we work out for illustration the leading constant $\mathfrak{S}(\mathbf{a},\Psi)$ in Theorem 1.2 in case $t=3$ for some values of $a$. For example, if $a=2$, we get

### v3

@@ Abstract.
We prove analogues of the theorem of Green and Tao on linear constellations in primes, in which the primes under consideration are restricted by certain arithmetic conditions. Our first main result is conditional upon Hooley’s Riemann hypothesis and imposes the extra condition that the primes have prescribed primitive roots. Our second main result is unconditional and imposes the extra condition that the primes have prescribed Artin symbols in given Galois number fields. In the appendix we present an application of the second result in inverse Galois theory.
@@ 1.1. Primes with prescribed primitive roots
simultaneously take values in the set of primes with primitive root $5$. This is a constellation of complexity $2$ in the sense of Green and Tao. Their celebrated result [14], together with [15, 16] and the work [17] of Green, Tao and Ziegler, gives asymptotics for all constellations of finite complexity in unrestricted primes, implying in particular the existence of infinitely many solutions $(n_{1},n_{2})$ to our question if one ignores the primitive root condition.
@@
Here, $\mu_{k}$ denotes the group of $k$-th roots of unity in $\mathbb{C}$. For background and history surrounding Artin’s conjecture and Hooley’s result, we recommend Moree’s survey [36]. Our first main result establishes, conditionally on $\operatorname{HRH}(7)$, the existence of infinitely many solutions $(n_{1},n_{2})$ as above. In fact, we allow arbitrary primitive roots, arbitrary finite-complexity systems of affine-linear forms, and obtain an asymptotic formula analogous to the theorem of Green and Tao for unrestricted primes.
@@
$\left\{p\leq N\ :\ p\equiv b\bmod q,\ \mathbb{F}_{p}^{\times}=\langle a\rangle\right\}=\delta(a,b,q)\frac{N}{\log N}+o_{a,q}\left(\frac{N}{\log N}\right).$ (1.2)
@@
[…]
@@
where we have used the notation $N]=\{1,\ldots,N\}$. Our main result extends this from $\psi(n)=b+qn$ to arbitrary finite complexity [^1] Recall the characterisation of finite complexity from [14, Lemma 1.6]: no two of the forms $\psi_{i}$ have linearly dependent linear parts. systems of affine-linear forms $\Psi=(\psi_{1},\ldots,\psi_{t}):\mathbb{Z}^{s}\to\mathbb{Z}^{t}$ as in [14, Definition 1.1 and Definition 1.5]. For $N\in\mathbb{N}$, the size of $\Psi$ relative to $N$ is defined as
@@
where $\dot{\psi}_{i}=\psi_{i}-\psi_{i}(0)$ is the linear part of $\psi_{i}$ and $e_{1},\ldots,e_{s}$ is the standard basis of $\mathbb{Z}^{s}$. Let $\mathbf{a}=(a_{1},\ldots,a_{t})\in\mathbb{Z}^{t}$ such that no $a_{i}$ is equal to 1-1 or a perfect square. For any $q\in\mathbb{N}$, we define the density
@@
$\mathscr{D}_{\mathbf{a}}:=\operatorname{lcm}(|\Delta_{a_{i}}|\ :\ 1\leq i\leq t).$
With this setup in place, we can state our first main result, a version of [14, Main Theorem] with prescribed primitive roots, conditional upon the minimal version of GRH that is currently needed to even establish the existence of infinitely many primes with these prescribed primitive roots.
@@
Let $s,t,L,N\in\mathbb{N}$. Let $\mathbf{a}=(a_{1},\ldots,a_{t})\in\mathbb{Z}^{t}$ such that no $a_{i}$ is equal to 1-1 or a perfect square, and assume that $\operatorname{HRH}(a_{i})$ holds for all $1\leq i\leq t$. Let $\Psi=(\psi_{1},\ldots,\psi_{t}):\mathbb{Z}^{s}\to\mathbb{Z}^{t}$ be a system of affine-linear forms of finite complexity with size $\lVert\Psi\rVert_{N}\leq L$. Let $X\subseteq[-N,N]^{s}$ be a convex set. Then
@@
for unrestricted primes in [14, (1.6)], where $\phi(q)$ denotes Euler’s totient and for any proposition $P$ we write $\mathbf{1}_{P}=1$ if $P$ is true and $\mathbf{1}_{P}=0$ otherwise. Here, $\mathbf{1}_{\gcd(\psi_{i}(n),q)=1}/\phi(q)$ can be seen as the probability that a random prime is congruent to $\psi_{i}(n)\bmod q$ by the prime number theorem in arithmetic progressions. Contrary to [14], our factor $\mathfrak{S}(\Psi,\mathbf{a})$ does not factor completely as an Euler product, hence the need to introduce the quantity $\mathscr{D}_{\mathbf{a}}$. Moreover, our formula features an additional global factor $\prod_{i=1}^{t}\delta(a_{i},0,1)$, where of course $\delta(a_{i},0,1)$ should be seen as the probability that a random prime has primitive root $a_{i}$. Both of these phenomena were already observed in a special case in [10], see Example 1.8 below.
@@
If $\sigma_{\mathbf{a},\Psi}(q)=0$ for some $q\in\{\mathscr{D}_{\mathbf{a}}\}\cup\{p\ :\ p\nmid\mathscr{D}_{\mathbf{a}},\ p\ll_{s,t,L,\mathbf{a}}1\}$, then for each $c\in(\mathbb{Z}/q\mathbb{Z})^{s}$ there is some $i\in[t]$ with $\delta(a_{i},\psi_{i}(c),q)=0$. From Moree’s explicit description of $\delta(a,b,q)$, see §3.3, it is straightforward to see that then every prime congruent to $\psi_{i}(c)\bmod q$ with primitive root $a_{i}$ must lie in the set $\{2,\gcd(\psi_{i}(c),q)\}$. Taking all possible candidates for all $c$, we end up with a finite collection of cases, in each of which one of the forms $\psi_{i}(n)$ has to take a fixed value. In each case, one can use linear algebra to either show that no solutions exist, or express the system in s1s-1 variables. However, this new system might have infinite complexity, so Theorem 1.2 might not apply any longer. See the end of §11.3 for an example.
@@ 1.2. Primes with prescribed Artin symbols
Our second main result is unconditional. Here, instead of prescribed primitive roots, we require the primes to have prescribed Artin symbols $K/\mathbb{Q},p]$ for some finite Galois extensions $K/\mathbb{Q}$. Recall that this Artin symbol is defined for primes $p$ unramified in $K$ and yields a conjugacy class in the Galois group $\operatorname{Gal}(K/\mathbb{Q})$, the class of Frobenius elements of prime ideals $\mathfrak{p}$ of the ring of integers of $K$ lying above $p$. For any such conjugacy class $C$, we define the Chebotarev-von Mangoldt function
[…]
@@
[…]
@@
$\mathscr{D}_{\mathbf{K}}:=\operatorname{lcm}(\Phi_{K_{i}^{\operatorname{ab}}}\ :\ 1\leq i\leq t),$ (1.14)
where $\Phi_{K_{i}^{\operatorname{ab}}}\in\mathbb{N}$ is the finite part of the conductor [^2] In concrete terms, $\Phi_{K_{i}^{\operatorname{ab}}}$ is the minimal $f\in\mathbb{N}$ such that $K_{i}^{\operatorname{ab}}\subseteq\mathbb{Q}(\mu_{f})$. Such $f$ exist by the Kronecker-Weber theorem. of the maximal abelian subextension $K_{i}^{\operatorname{ab}}/\mathbb{Q}$ of $K_{i}/\mathbb{Q}$. With this setup in place, our second main result is the following version of the theorem of Green and Tao.
@@
The factor $\mathfrak{S}(\mathbf{K},\mathbf{C},\Psi)$ in the main term has an analogous interpretation as in Theorem 1.2. In particular, the quotient $\eta_{K_{i},C_{i}}(\psi_{i}(n),q)/\eta_{K_{i},C_{i}}(0,1)$ in (1.13) can be seen as the conditional probability that a random prime $p$ with Artin symbol $K_{i}/\mathbb{Q},p]=C_{i}$ is congruent to $\psi_{i}(n)$ modulo $q$. In contrast to Theorem 1.2, these conditional probabilities simplify considerably: we will see in Lemma 3.4 that $\tau_{\mathbf{K},\mathbf{C},\Psi}(\mathscr{D}_{\mathbf{K}})=\beta_{\Psi,p}$ whenever $p\nmid\mathscr{D}_{K}$, with $\beta_{\Psi,p}$ the same density as for unrestricted primes in [14], defined in (1.8). Hence, we can also write
@@
which is very close to the formula for unrestricted primes in [14, Main Theorem].
@@
Analogous statements as in Remark 1.3 apply. One easily sees directly from the definition that if $\eta_{K,C}(b,q)=0$, then there is at most one possible prime $p\equiv b\bmod q$ with $K/\mathbb{Q},p]=C$, namely $\gcd(b,q)$.
@@
The theorem of Green and Tao [14] has been applied in combination with fibration or specialisation techniques in more algabraically-flavored contexts, where it is used to show that certain linear forms, e.g. the linear factors of a well-chosen discriminant polynomial, take prime values simultaneously infinitely often (see e.g. [18, 3, 25, 1]). Theorem 1.4 adds the capability of prescribing splitting conditions at these prime values, which we hope will be useful. To facilitate applications, we state the following implication of Theorem 1.4.
@@
Then there are infinitely many $n\in\mathbb{Z}^{s}\cap X$ for which the values $\psi_{i}(n)$ are distinct primes with $K_{i}/\mathbb{Q},\psi_{i}(n)]=C_{i}$ for all $i$.
@@ 1.3. Application in inverse Galois theory
A special case of Corollary 1.6 was asked for explicitly by Kim and König in the context of inverse Galois problems with local conditions. In [25, Remark 6], they ask whether, for given $n\geq 2$, an analog of the theorem of Green and Tao can prove the existence of infinitely many pairs $(s,t)\in\mathbb{Z}^{2}$, such that
$|t|$ $\text{ is a prime congruent to }1\bmod n,$
$|s|$ $\text{ is a prime congruent to }1\bmod n-1,$
$|(n-1)^{n-1}t-n^{n}s|$ $\text{ is a prime splitting completely in }F_{n-1},$
@@
over $\mathbb{Q}$. In Appendix A, which is joint work with Joachim König, we apply Corollary 1.6 to answer this question positively for all $n\geq 2$. Using this to replace an application of the original theorem of Green and Tao in the proof of [25, Theorem 5.5], we prove the following result on locally cyclic $\operatorname{S}_{n}$-extensions with prescribed Artin symbol at finitely many primes. This generalisation of the special case $n=5$ treated in [25, Theorem 5.5] was suggested by Kim and König as the motivation for their question. Recall that a Galois number field $K/\mathbb{Q}$ is locally cyclic if all decomposition groups are cyclic.
@@
    For each $p\in M_{2}$, we have $K/\mathbb{Q},p]=C_{p}$.
This result is new even without the conditions (3) and (4). It answers positively a question of Bubboloni and Sonn [8], who asked whether locally cyclic Galois extensions of $\mathbb{Q}$ with Galois group $\operatorname{S}_{n}$ exist for all $n$. Let us note that [25, Theorem 5.5] also covers the case of $\operatorname{PGL}_{2}(7)$. A generalisation to $\operatorname{PGL}_{2}(p)$ for arbitrary primes $p$ follows from the main result of [1].
@@ 1.4. Previous results and examples
Actually, Kane’s formula is stated in terms of primes $p_{i}$ with $K_{i}/\mathbb{Q},p_{i}]=C$ weighted by $\log(p_{i})$ instead of integers $n_{i}$ weighted by $\Lambda_{K_{i},C_{i}}(n_{i})$, but the contribution of prime powers making up the difference between these two settings is clearly negligible. Considering $c_{1},\ldots,c_{t}$ as fixed, we may assume that $M\ll N$, as otherwise the sum is zero. As explained in [14, Theorem 1.8] and its proof, one may choose a basis for the affine sublattice of $\mathbb{Z}^{t}$ given by $c_{1}x_{1}+\cdots+c_{t}x_{t}=M$, and thus describe it as the image of an affine-linear system $\Psi:\mathbb{Z}^{t-1}\to\mathbb{Z}^{t}$ of size $\lVert\Psi\rVert\ll 1$ and complexity $\leq 1$.
We describe Kane’s formula in §11.1 and show that it agrees with the corresponding special case of Theorem 1.4. In order to obtain his result, Kane generalises the proof of Vinogradov’s three primes theorem presented in [21, §19], based on classical analytic number theory surrounding the Hardy-Littlewood circle method. Some of Kane’s techniques play a role in our work. Most importantly, our Proposition 5.2 can be seen as a higher-order Fourier analysis version of [24, Lemma 21], Kane’s exponential sum estimate for polynomial phases evaluated at ideal norms.
We should mention that Kane’s result features a quantitative error term saving an arbitrary power of $\log N$, whereas our error terms are purely qualitative. There has been great recent progress in obtaining quantitative versions of the methods of Green, Tao and Ziegler. Tao and Teräväinen [40], based on Manners’ quantitative inverse theorem for Gowers norms [31], obtained a version of the theorem of Green and Tao with quantitative error terms saving a small power of $\log\log N$. A very recent breakthrough in the area is the quasipolynomial inverse theorem of Leng, Sah and Sawhney [29]. One might hope that applications of this work could lead to quantitative error terms and also dispose of the need for pseudorandom majorants in our work. In the context of Theorem 1.2, this does not seem very promising with what we have at the moment, as our bounds for correlations with nilsequences are extremely inexplicit, see Theorem 2.6. Although we are not following this direction in the present work, we would be very interested in seeing explicit bounds in the context of Theorem 1.4.
@@
of complexity t2t-2 takes infinitely many values in $\mathscr{P}^{t}$. Taking $\mathscr{P}$ to be the set of primes with $K/\mathbb{Q},p]=C$ for some fixed $K,C$, which has positive density in the primes by (1.12) with $q=1$, one obtains a weakened special case of Theorem 1.4. Taking $\mathscr{P}$ to be the set of primes with primitive root $a$, one obtains a weakened special case of Theorem 1.2. In both cases, the precise asymptotic formulae in our results are new. In §11.3, we work out for illustration the leading constant $\mathfrak{S}(\mathbf{a},\Psi)$ in Theorem 1.2 in case $t=3$ for some values of $a$. For example, if $a=2$, we get